# Skyfall-GS: Synthesizing Immersive 3D Urban Scenes from Satellite Imagery

### v1

@@ Abstract
Synthesizing large-scale, explorable, and geometrically accurate 3D urban scenes is a challenging yet valuable task in providing immersive and embodied applications. The challenges lie in the lack of large-scale and high-quality real-world 3D scans for training generalizable generative models. In this paper, we take an alternative route to create large-scale 3D scenes by synergizing the readily available satellite imagery that supplies realistic coarse geometry and the open-domain diffusion model for creating high-quality close-up appearances. We propose Skyfall-GS, the first city-block scale 3D scene creation framework without costly 3D annotations, also featuring real-time, immersive 3D exploration. We tailor a curriculum-driven iterative refinement strategy to progressively enhance geometric completeness and photorealistic textures. Extensive experiments demonstrate that Skyfall-GS provides improved cross-view consistent geometry and more realistic textures compared to state-of-the-art approaches. Project page: https://skyfall-gs.jayinnn.dev/.
@@ 1 Introduction
Synthetic high-quality, immersive, and semantically plausible 3D urban scenes have crucial applications in gaming, filmmaking, and robotics. The ability to create a large-scale and 3D-grounded environment supports realistic rendering and immersive experience for storytelling, demonstration, and embodied physics simulation. However, due to limited 3D-informed data, building a generative model for realistic and navigable 3D cities remains challenging. It is expensive and labor-intensive to acquire large-scale 3D and textured reconstructions of cities with detailed geometry, while using Internet image collections faces challenges in camera pose registration and excessive data noise (e.g., transient objects and different times of the day). These constraints set back existing 3D city generation frameworks from creating realistic and diverse appearances. With this observation, we propose an alternative route for virtual city creation with a two-stage pipeline: partial and coarse geometry reconstruction from multi-view satellite imagery, then close-up appearance completion and hallucination using an open-domain diffusion model.
@@
Satellite imagery offers a compelling alternative due to its extensive geographic coverage, automated collection, and high-resolution capabilities. For instance, Maxar’s WorldView-3 satellite captures approximately 680,000 $\text{km}^{\text{2}}$ of imagery daily at resolutions up to 31 cm per pixel. Such data inherently encodes semantically plausible representations of real-world environments, enabling scalable 3D urban scene creation. However, in Figure˜2(a), we show that directly applying 3D reconstruction methods to satellite imagery is insufficient for creating navigable and immersive 3D cities. The significant amount of invisible regions (e.g., building facades) and limited satellite-view parallax create incorrect geometry and artifacts.
@@
Completing and enhancing the geometry and texture in the ground view requires a significant influx of extra information. In Figure˜2(b), we study a few state-of-the-art methods in city generation (Xie et al., 2024; 2025b). These methods produce oversimplified building geometries and unrealistic appearances due to strong assumptions, particularly the reliance on semantic maps and height fields as the sole inputs, and overfitting to small-scale, domain-specific datasets. Such an observation motivates us to seek help from open-domain foundation vision models as an external information source, which provides better zero-shot generalization and diversity. Noticing that the ground-view novel-view renderings from the GS reconstructed scene exhibit noise-like patterns, we treat these renderings as intermediate results in a denoising diffusion process. Then, we complete the remaining denoising process to create hallucinated pseudo ground-truth for the GS scene optimization. To stabilize the convergence, we carefully design a curriculum-based view selection and iterative refinement process, where the sampled view angles gradually fall from the sky to the ground over time. Accordingly, we name our framework Skyfall-GS. In Figure˜1 and Figure˜2, we show that Skyfall-GS yields significantly enhanced texture with 3D-justified geometry compared to the relevant baselines.
@@
Skyfall-GS is the first method synthesizing immersive, free-flight navigable 3D urban scenes without fixed-domain training on 3D data. Skyfall-GS operates on readily available satellite imagery as the only input, then hallucinates realistic aerial-view appearances and maintains a strong satellite-to-ground 3D consistency. Moreover, Skyfall-GS supports real-time and interactive rendering, as we design our framework to produce GS results without sophisticated data structures. Through experiments on diverse environments, we show that Skyfall-GS has better generalization and robustness compared to state-of-the-art methods. Our ablation shows that each of our designs improves the perceptual plausibility and semantic consistency. Skyfall-GS paves the way for scalable 3D urban virtual scene creation, enabling applications in virtual entertainment, simulation, and robotics.
@@
  • We introduce Skyfall-GS, the first method to synthesize immersive, real-time free-flight navigable 3D urban scenes solely from multi-view satellite imagery using generative refinement.
  • An open-domain refinement approach leveraging pre-trained text-to-image diffusion models without domain-specific training.
  • A curriculum-learning-based iterative refinement strategy progressively enhances reconstruction quality from higher to lower viewpoints, significantly improving visual fidelity in occluded areas.
Figure 2: Limitations of existing novel-view synthesis methods from satellite imagery. (a) Sat-NeRF (Marí et al., 2022) and naive 3DGS (Kerbl et al., 2023) yield blurred or distorted building facades due to insufficient geometric detail and limited parallax from satellite viewpoints. (b) City generation methods (Xie et al., 2024; 2025b) produce oversimplified building geometries and unrealistic appearances, primarily due to strong assumptions about the input data, and overfitting to small-scale, domain-specific datasets. In comparison, our method synthesizes more realistic appearances and geometries from aerial views.
@@ 2 Related Work
Gaussian Splatting. 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) offers real-time view synthesis rivaling NeRFs (Mildenhall et al., 2021; Barron et al., 2021; 2022; Müller et al., 2022; Barron et al., 2023; Martin-Brualla et al., 2021). Mip-Splatting (Yu et al., 2024) fixes scale-change issues via on-the-fly resizing. Recent advances target satellite and aerial reconstruction: FusionRF (Sprintson et al., 2024) achieves 17% depth improvement from multispectral acquisitions, while InstantSplat (Fan et al., 2024) enables 40-second pose-free reconstruction. “In-the-wild” variants handle appearance and uncertainty (Xu et al., 2024; Sabour et al., 2024; Wang et al., 2024b; Dahmani et al., 2024; Zhang et al., 2024a; Kulhanek et al., 2024; Hou et al., 2025), including SpectroMotion (Fan et al., 2025) for dynamic specular scenes, while large-scene methods use LOD and partitioning (Kerbl et al., 2024; Liu et al., 2025c; 2024; Lin et al., 2024; Turki et al., 2022; Tancik et al., 2022). CAT-3DGS (Zhan et al., 2025) achieves rate-distortion optimization via context-adaptive triplanes. For sparse-view satellite imagery, depth or co-regularization priors guide reconstruction (Li et al., 2024b; Zhang et al., 2024b; Zhu et al., 2023; Niemeyer et al., 2022; Lin et al., 2025), with SparseSat-NeRF (Zhang & Rupnik, 2023) adding dense depth supervision.
Diffusion models for 3D reconstruction and editing. Diffusion models (Rombach et al., 2022; Labs, 2024b) underpin image generation and editing. Early SDS pipelines DreamFusion (Poole et al., 2022) and Magic3D (Lin et al., 2023a) enabled text-to-3D, with ProlificDreamer (Wang et al., 2023) addressing over-smoothing via Variational Score Distillation. DreamGaussian (Tang et al., 2023) achieves 10x speedup via progressive densification, while GaussianDreamer (Yi et al., 2024) bridges 2D and 3D diffusion models. SDEdit (Meng et al., 2022), DDIM inversion (Mokady et al., 2022; Miyake et al., 2024), and FlowEdit (Kulikov et al., 2024) enable fine control. Extensions include sparse-view reconstruction (Wu et al., 2023; Liu et al., 2023b; Chen et al., 2024), with MVDream (Shi et al., 2023) enabling multi-view consistency. For 3D/4D generation (Gao et al., 2024b; Wu et al., 2024b; Melas-Kyriazi et al., 2024; Chung et al., 2023; Liu et al., 2023a) and scene editing (Haque et al., 2023; Wu et al., 2025; Ye et al., 2024b; Fang et al., 2024; Mirzaei et al., 2024; Dihlmann et al., 2024; Weber et al., 2024; Wu et al., 2024a; Wang et al., 2025), SPIn-NeRF (Mirzaei et al., 2023) handles occlusions via perceptual inpainting while CF-NeRF (Shen et al., 2022) provides uncertainty quantification. CorrFill (Liu et al., 2025a) enhances faithfulness via correspondence guidance, while AuraFusion360 (Wu et al., 2025) enables 360° scene inpainting for Gaussian Splatting. Instruct-NeRF2NeRF (Haque et al., 2023) refines NeRF views iteratively with Instruct-Pix2Pix (Brooks et al., 2023) for diffusion-driven 3D editing.
Urban scene modeling. Classic SfM-MVS pipelines extract DSMs from satellite pairs (Schönberger & Frahm, 2016; Zhang et al., 2019; Gao et al., 2023a), with MVS3D (Bosch et al., 2016) benchmarks for evaluation. Neural variants improve geometric fidelity (Derksen & Izzo, 2021a; Marí et al., 2022; 2023; Zhou et al., 2024b; Leotta et al., 2019; Liu et al., 2025b; Qu & Deng, 2023; Gao et al., 2024a; Savant Aira et al., 2025; Huang et al., 2025), including Planet-NeRF (Derksen & Izzo, 2021b) for global-scale scenes and SatMVS (Gao et al., 2021; 2023b) with RPC warping, yet both miss occluded facades. Generative synthesis divides into: (i) street-view methods (Li et al., 2024c; 2021; d; Toker et al., 2021; Qian et al., 2023; Shi et al., 2022; Ze et al., 2025; Deng et al., 2024; Xu & Qin, 2025), including GeoDiffusion (Xiong et al., 2024) for mixed-view synthesis, Geospecific View Generation (Xu & Qin, 2024) achieving 10x resolution gains, and SkyDiffusion (Ye et al., 2024a) with Curved-BEV for street-to-satellite mapping, though lacking 3D consistency and temporal coherence; and (ii) full-3D city generation (Lin et al., 2023b; Xie et al., 2024; 2025a; 2025b; Sun et al., 2024; Zhou et al., 2024a; Shang et al., 2024; Li et al., 2024a; Zhang et al., 2024c), with BEVFormer (Li et al., 2022) and MagicDrive (Gao et al., 2023c) using spatiotemporal transformers for view consistency. While Infinicity (Lin et al., 2023b) uses pixel-to-voxel rendering for infinite cities, and CityDreamer (Xie et al., 2024) and GaussianCity (Xie et al., 2025b) use BEV neural fields or BEV-Point splats for editable scenes, these remain constrained by input representations (semantic maps and height fields) and training distributions, limiting synthesis of realistic textures and complex structures like tunnels, bridges, and multi-level architectures. Our method uses pretrained diffusion priors to recover high-fidelity facades in occluded regions without dataset-specific training, respecting user constraints more faithfully.
@@ 3 Method
Figure 3: Overview of the proposed Skyfall-GS pipeline. Our method synthesizes immersive and free-flight navigable city-block scale 3D scenes solely from multi-view satellite imagery in two stages. (a) In the Reconstruction Stage, we first reconstruct the initial 3D scene using 3DGS, enhanced by pseudo-camera depth supervision to address limited parallax in satellite images. We use an appearance modeling component to handle varying illumination conditions across multi-date satellite images. (b) In the Synthesis Stage, we introduce a curriculum-based Iterative Dataset Update (IDU) refinement technique leveraging (c) a pre-trained T2I diffusion model (Labs, 2024b) with prompt-to-prompt editing (Kulikov et al., 2024). By iteratively updating training datasets with progressively refined renders, our approach significantly reduces visual artifacts, improving geometric accuracy and texture realism, particularly in previously occluded areas such as building facades.
Our two-stage pipeline (Figure˜3) turns satellite images into immersive 3D cities. Reconstruction Stage (Section˜3.1): fit a 3D Gaussian Splatting model, adding illumination-adaptive appearance modeling and regularizers for sparse, multi-date views. Synthesis Stage (Section˜3.2): recover occluded regions, e.g., facades, through curriculum Iterative Dataset Update, repeatedly refining renders with text-guided diffusion edits. The loop keeps textures faithful to the satellite input while preserving geometry, yielding complete, navigable urban scenes from satellite data alone.
Preliminary. 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) encodes a scene as Gaussians with center $\mu_{i}$, covariance $\Sigma_{i}$, opacity $\alpha_{i}$, and view-dependent color. Each Gaussian projects to the image plane with covariance: $\Sigma_{i}^{\prime}=JW\Sigma_{i}W^{T}J^{T}$, where $W$ is the viewing transformation and $J$ is the affine-projection Jacobian. Pixels are alpha-composited front-to-back. Parameters are trained with:
$\mathcal{L}_{\text{color}}=\lambda_{\text{D-SSIM}}\,\text{DSSIM}(\hat{C},C)+(1-\lambda_{\text{D-SSIM}})\lVert\hat{C}-C\rVert_{1}\,\,.$ (1)
@@ 3.1 Initial 3DGS Reconstruction from Satellite Imagery
Approximated camera parameters. Satellite imagery typically uses the rational polynomial camera (RPC) model, directly mapping image coordinates to geographic coordinates. To integrate with the 3DGS pipeline, we employ SatelliteSfM (Zhang et al., 2019) to approximate perspective camera parameters (extrinsic and intrinsic) from RPC and generate sparse SfM points as initial 3DGS points.
Appearance modeling. As highlighted in Section˜1, multi-date satellite imagery exhibits significant appearance variations due to global illumination changes, seasonal factors, and transient objects, as illustrated in Figure˜3(a). Following WildGaussians (Kulhanek et al., 2024), we use trainable per-image embeddings $\{e_{j}\}_{j=1}^{N}$ (with $N$ training images) to handle varying illumination and atmospheric conditions. We also employ trainable per-Gaussian embeddings $g_{i}$ to capture localized appearance changes like shadow variations. A lightweight MLP $f$ computes affine color transformation parameters $(\beta,\gamma)$ as $(\beta,\gamma)=f(e_{j},g_{i},\bar{c}_{i}),$ where $e_{j}$ is the per-image embedding, $g_{i}$ is the per-Gaussian embedding, and $\bar{c}_{i}$ denotes the 0-th order spherical harmonics (SH). Finally, the transformed color $\tilde{c}_{i}$ is then computed as $\tilde{c}_{i}(\mathbf{r})=\gamma\cdot\hat{c}_{i}(\mathbf{r})+\beta,$ and used in the 3DGS rasterizer. To prevent modeling the appearance changes as view-dependent effects, we limit SH to zero and first-order terms.
Opacity regularization. We observed that numerous floaters in reconstructed scenes exhibit low opacity. To encourage geometry to adhere closely to actual surfaces, we propose entropy-based opacity regularization:
$\mathcal{L}_{\text{op}}=-\sum_{i}\alpha_{i}\log(\alpha_{i})+(1-\alpha_{i})\log(1-\alpha_{i})\,\,.$ (2)
This regularization promotes binary opacity distributions, allowing low-opacity Gaussians to be pruned during densification. Incorporating this term significantly sharpens geometric reconstruction, providing a better foundation for subsequent synthesis.
Pseudo camera depth supervision. To further reduce floating artifacts, we sample pseudo-cameras positioned closer to the ground during optimization. From these pseudo-cameras, we render RGB images $I_{\text{RGB}}$ and corresponding alpha-blended depth maps $\hat{D}_{\text{GS}}$. We then use an off-the-shelf monocular depth estimator, MoGe (Wang et al., 2024a), to predict scale-invariant depths $\hat{D}_{\text{est}}$ from these renders. We use the absolute value of Pearson correlation (PCorr) to supervise the depth:
$\mathcal{L}_{\text{depth}}=\lVert\text{PCorr}(\hat{D}_{\text{GS}},\hat{D}_{\text{est}})\rVert_{1}\,\,;\quad\text{PCorr}(\hat{D}_{\text{GS}},\hat{D}_{\text{est}})=\frac{\text{Cov}(\hat{D}_{\text{GS}},\hat{D}_{\text{est}})}{\sqrt{\text{Var}(\hat{D}_{\text{GS}})\text{Var}(\hat{D}_{\text{est}})}}\,\,.$ (3)
Optimization. Combining all components, the overall loss for the reconstruction stage is defined as:
$\mathcal{L}_{\text{sat}}(G,C)=\mathcal{L}_{\text{color}}+\lambda_{\text{op}}\mathcal{L}_{\text{op}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}\,\,$, (4)
@@
3.2 Synthesize via Curriculum-learning Based Iterative Datasets Update
The iterative dataset update (IDU) technique (Haque et al., 2023; Melas-Kyriazi et al., 2024) repeatedly executes render-edit-update cycles across multiple episodes to progressively synthesize 3D scenes. Unlike previous methods that sample camera poses from original training views (Haque et al., 2023) or simple orbits (Melas-Kyriazi et al., 2024), we introduce a curriculum-based refinement schedule over $N_{e}$ episodes that specifically addresses satellite imagery’s geometric and visual limitations, producing structurally accurate and photorealistic reconstructions of occluded areas.
Curriculum learning strategy. As illustrated in Figure˜5, we observe that 3DGS trained from satellite imagery produces higher-quality renders at higher elevation angles but degenerates at lower elevation angles. Leveraging this insight, we introduce a curriculum-based synthesizing strategy, which progressively lowering viewpoints across optimization episodes. Specifically, we define $N_{p}$ look-at points $\{P_{i}\}_{i=1}^{N_{p}}$ uniformly placed throughout the scene and uniformly sample $N_{v}$ camera positions along orbital trajectories with controlled elevation angles and radii. Our iterative dataset update (IDU) process starts from higher elevations, progressively moving toward lower perspectives. This approach gradually reveals previously occluded regions, improving geometric detail and texture realism, as validated in our ablation studies (Section˜4.2).
Render refinement by text-to-image diffusion model. As illustrated in Figure˜5(a), renderings from initial 3DGS contain blurry texture and artifacts. To address this, we leverage prompt-to-prompt editing with pre-trained text-to-image diffusion models to synthesize disocclusion areas, remove artifacts, and enhance geometry. Prompt-to-prompt editing (Hertz et al., 2022) modifies input images, which are described by the source prompt, to align with the target prompt while preserving structural content. Although typically used on real or diffusion-generated photos, we demonstrate its effectiveness for refining degraded satellite-trained 3DGS renders. We employ FlowEdit (Kulikov et al., 2024) with the pre-trained FLUX.1 [dev] diffusion model (Labs, 2024a), using prompt pairs that transform degenerate renders into high-quality imagery. Our prompts specifically describe the degraded features in original renders and specify the desired high-quality attributes in target prompts, see Section˜A.1 for prompts detail. As illustrated in Figure˜5, this approach significant improves the visual quality of renders, including sharper geometric details, enhanced texture richness, and physically coherent shadows, strengthening the 3DGS training dataset for more accurate reconstructions.

### v2

@@ Abstract
Synthesizing large-scale, explorable, and geometrically accurate 3D urban scenes is a challenging yet valuable task for immersive and embodied applications. The challenge lies in the lack of large-scale and high-quality real-world 3D scans for training generalizable generative models. In this paper, we take an alternative route to create large-scale 3D scenes by leveraging readily available satellite imagery for realistic coarse geometry and open-domain diffusion models for high-quality close-up appearance synthesis. We propose Skyfall-GS, a novel hybrid framework that synthesizes immersive city-block scale 3D urban scenes by combining satellite reconstruction with diffusion refinement, eliminating the need for costly 3D annotations, and also featuring real-time, immersive 3D exploration. We tailor a curriculum-driven iterative refinement strategy to progressively enhance geometric completeness and photorealistic texture. Extensive experiments demonstrate that Skyfall-GS provides improved cross-view consistent geometry and more realistic textures compared to state-of-the-art approaches. Project page: https://skyfall-gs.jayinnn.dev/
@@ 1 Introduction
High-quality, immersive, and semantically plausible 3D urban scenes are essential for a wide range of applications, including gaming, filmmaking, navigation planning, and robotics. The ability to create a large-scale and 3D-grounded environment supports realistic rendering and immersive experience for storytelling, demonstration, and embodied physics simulation. However, due to limited 3D-informed data, building a generative model for realistic and navigable 3D cities remains challenging. It is expensive and labor-intensive to acquire large-scale 3D and textured reconstructions of cities with detailed geometry, while using Internet image collections faces challenges in camera pose registration and excessive data noise (e.g., transient objects and different times of the day). These constraints prevent existing 3D city generation frameworks from creating realistic and diverse appearances. With this observation, we propose an alternative route for virtual city creation with a two-stage pipeline: partial and coarse geometry reconstruction from multi-view satellite imagery, then close-up appearance completion and synthesis using an open-domain diffusion model.
@@
Satellite imagery offers a compelling alternative due to its extensive geographic coverage, automated collection, and high-resolution capabilities. For instance, Maxar’s WorldView-3 satellite captures approximately 680,000 $\text{km}^{\text{2}}$ of imagery daily at resolutions up to 31 cm per pixel. Such data encodes a large volume of real-world environment semantics, enabling scalable 3D urban scene creation. However, in Figure˜2(a), we show that directly applying 3D reconstruction methods to satellite imagery is insufficient for creating navigable and immersive 3D cities. The substantial invisible regions (e.g., building facades) and limited satellite-view parallax create incorrect geometry and artifacts.
@@
Completing and enhancing the geometry and texture in the ground view requires a significant influx of extra information. In Figure˜2(b), we study a few state-of-the-art methods in city generation [101, 103]. These methods produce oversimplified building geometries and unrealistic appearances due to strong assumptions, particularly the reliance on semantic maps and height fields as the sole inputs, and overfitting to small-scale, domain-specific datasets. Such an observation motivates us to leverage an open-domain foundation image generation model as an external information source, providing better zero-shot generalization and diversity. Ground-level novel-view renderings from the GS-reconstructed scene suffer from severe degradation, including floater artifacts and texture smearing, caused by insufficiently constrained Gaussians in regions with limited satellite-view parallax. To address this, we employ an open-domain foundation image generation model to directly refine these degenerate renderings into photorealistic and geometrically consistent outputs, exploiting the model’s rich visual priors to recover plausible appearances. The refined outputs serve as pseudo ground-truth to supervise iterative GS scene optimizations, progressively improving the visual fidelity and geometric consistency across the scene. To stabilize the convergence, we carefully design a curriculum-based view selection and iterative refinement process, where the sampled view angles gradually fall from the sky to the ground over time. Accordingly, we name our framework Skyfall-GS. In Figure˜1 and Figure˜2, we show that Skyfall-GS substantially enhances texture with 3D-justified geometry compared to the relevant baselines.
@@
Skyfall-GS is a novel hybrid framework that synthesizes immersive 3D urban scenes by combining satellite reconstruction with diffusion refinement, eliminating the need for fixed-domain training on 3D data. Skyfall-GS operates on readily available satellite imagery as the only input, then synthesizes realistic aerial-view appearances and maintains a strong satellite-to-ground 3D consistency. Moreover, Skyfall-GS supports real-time and interactive rendering. Through experiments on diverse environments, we show that Skyfall-GS has better generalization and robustness compared to state-of-the-art methods. Our ablation study shows that each component improves perceptual plausibility and semantic consistency. Skyfall-GS paves the way for scalable 3D urban virtual scene creation, enabling applications in virtual entertainment, simulation, and robotics.
@@
  • We introduce Skyfall-GS, the first method to synthesize immersive, real-time free-flight navigable 3D urban scenes solely from multi-view satellite imagery using generative refinement without domain-specific training.
  • An open-domain generative refinement approach that exploits rich visual priors from pre-trained text-to-image diffusion models to recover photorealistic appearances from degenerate satellite reconstructions.
  • A curriculum-learning-based iterative refinement strategy that progressively enhances reconstruction quality from higher to lower viewpoints, significantly improving visual fidelity in occluded regions.
Figure 2: Limitations of existing novel-view synthesis methods from satellite imagery. (a) Sat-NeRF [56] and naive 3DGS [33] yield blurred or distorted building facades due to insufficient geometric detail and limited parallax from satellite viewpoints. (b) City generation methods [101, 103] produce oversimplified building geometries and unrealistic appearances, primarily due to strong assumptions about the input data, and overfitting to small-scale, domain-specific datasets. In comparison, our method synthesizes more realistic appearances and geometries from aerial views.
@@ 2 Related Work
2.0.1 Gaussian Splatting.
3D Gaussian Splatting (3DGS) [33] offers real-time view synthesis rivaling NeRFs [61, 1, 2, 66, 3, 58], with Mip-Splatting [111] addressing scale-change aliasing. In-the-wild variants handle appearance variation and transient objects [104, 76, 92, 10, 113, 35, 26, 14], while scaling to large scenes is addressed via partitioning and LOD [34, 49, 89, 86, 8, 21]. For sparse-view settings, depth and co-regularization priors guide reconstruction [42, 114, 121, 67, 48, 116].
2.0.2 Satellite and aerial 3D reconstruction.
Classical SfM-MVS pipelines extract DSMs from satellite imagery [78, 115, 17, 5], with neural variants improving geometric fidelity [12, 57, 120, 41, 52, 72, 19]. Sat-NeRF [56] and SatMVS [16, 18] apply NeRF and RPC-based warping to satellite imagery respectively, yet neither recovers occluded facades, while FusionRF [84] improves depth via multispectral acquisitions. EOGS [77] adapts 3DGS with affine cameras and shadow mapping for sparse multi-date satellite imagery, and SkySplat [28] outputs Gaussians from satellite imagery in a feed-forward manner. For aerial imagery, AGS [100] introduces Ray-Gaussian Intersection for large-scale surface reconstruction, CityGaussian [54, 55] achieves city-scale rendering via scene partitioning and LOD, and Horizon-GS [31] unifies aerial-to-ground reconstruction. However, no existing method jointly handles multi-date appearance variations in satellite imagery while synthesizing ground-level perspectives.
2.0.3 Urban scene synthesis.
Cross-view synthesis methods [74, 82, 105, 106, 112, 11, 88] generate ground-level images or videos from satellite imagery without explicit 3D representations, while others employ intermediate 3D structures such as point clouds [45], density fields [71, 69], or voxels [43], and recent methods [44, 107, 70, 27, 32] advance toward explicit 3D generation via diffusion- and lifting-based approaches. Layout-conditioned methods synthesize cities from BEV maps: InfiniCity [47] lifts infinite-pixel BEV maps to 3D via octree voxel completion and neural rendering, CityDreamer [101] introduces compositional generative models separating buildings from backgrounds, and GaussianCity [103] extends this paradigm to 3D Gaussian Splatting [102]; LLM-driven procedural approaches [118, 119, 79] further generate layouts from text, OSM, or semantic maps. Although CityX [118] can accept satellite images as layout guidance, none of these methods faithfully synthesize photorealistic ground-level views directly conditioned on the observed satellite texture.
2.0.4 Diffusion models for 3D reconstruction and editing.
Diffusion models [75, 38] have emerged as powerful generative priors for image synthesis. Score distillation pipelines [68, 46, 93] lift 2D priors into 3D; DreamGaussian [87] and GaussianDreamer [109] extend this to Gaussian Splatting, while MVDream [81] and sparse-view methods [99, 53, 7, 20, 98, 59, 9, 51] address multi-view consistency. Inversion-based methods [60, 65, 64, 36] and occlusion-aware inpainting works [63, 50, 96, 108, 15, 62, 13, 95, 97, 91, 80] enable scene editing and 3D Gaussian inpainting. Instruct-NeRF2NeRF [23] introduced the Iterative Dataset Update (IDU) paradigm, iteratively refining a NeRF via InstructPix2Pix [6] edits; LucidDreamer [9] and WonderWorld [110] adopt IDU for progressive scene expansion via extrapolation, and RealmDreamer [83] generates forward-facing scenes via iterative inpainting and depth diffusion. However, none of these can be naively applied to our setting: inpainting-based approaches treat facade regions as unknown and synthesize them freely, whereas satellite imagery does capture building facades at an oblique angle, providing appearance constraints that must be respected. Ignoring this leads to synthesized facades inconsistent with the satellite observations.
@@ 3 Method
Figure 3: Overview of the proposed Skyfall-GS pipeline. Our method synthesizes immersive and free-flight navigable city-block scale 3D scenes solely from multi-view satellite imagery in two stages. (a) In the Reconstruction Stage, we first reconstruct the initial 3D scene using 3DGS, enhanced by pseudo-camera depth supervision to address limited parallax in satellite images. We integrate an appearance modeling to handle varying illumination conditions across multi-date satellite images. (b) In the Synthesis Stage, we introduce a curriculum-based Iterative Dataset Update (IDU) refinement technique leveraging (c) a pre-trained T2I diffusion model [38] with prompt-to-prompt editing [36]. By iteratively updating training datasets with progressively refined renders, our approach significantly reduces visual artifacts, improving geometric accuracy and texture realism, particularly in previously occluded areas such as building facades.
Our two-stage pipeline (Figure˜3) transforms satellite images into immersive 3D cities. In the Reconstruction Stage (Section˜3.1), we fit a 3D Gaussian Splatting model, adding illumination-adaptive appearance modeling and regularizers for sparse, multi-date views. In the Synthesis Stage (Section˜3.2), we recover occluded regions, e.g., facades, through curriculum-based Iterative Dataset Update, repeatedly refining renders with text-guided diffusion edits. The loop keeps textures faithful to the satellite input while preserving geometry, yielding complete, navigable urban scenes from satellite data alone.
3.0.1 Preliminary.
3D Gaussian Splatting (3DGS) [33] encodes a scene as Gaussians with center $\mu_{i}$, covariance $\Sigma_{i}$, opacity $\alpha_{i}$, and view-dependent color. Each Gaussian projects to the image plane with covariance: $\Sigma_{i}^{\prime}=JW\Sigma_{i}W^{T}J^{T}$, where $W$ is the viewing transformation and $J$ is the affine-projection Jacobian. Pixels are alpha-composited front-to-back. Parameters are trained with:
$\small\mathcal{L}_{\text{color}}=\lambda_{\text{D-SSIM}}\,\text{DSSIM}(\hat{C},C)+(1-\lambda_{\text{D-SSIM}})\lVert\hat{C}-C\rVert_{1}\,\,.$ (1)
@@ 3.1 Initial 3DGS Reconstruction from Satellite Imagery
3.1.1 Approximated camera parameters.
Satellite imagery typically uses the rational polynomial camera (RPC) model, directly mapping image coordinates to geographic coordinates. To integrate with the 3DGS pipeline, we employ SatelliteSfM [115] to approximate perspective camera parameters (extrinsic and intrinsic) from RPC and generate sparse SfM points as initial 3DGS points.
3.1.2 Appearance modeling.
[…]
3.1.3 Opacity regularization.
We observe that numerous floaters in reconstructed scenes exhibit low opacity. To encourage geometry to adhere closely to actual surfaces, we propose entropy-based opacity regularization:
$\small\mathcal{L}_{\text{op}}=-\sum_{i}\left[\alpha_{i}\log(\alpha_{i})+(1-\alpha_{i})\log(1-\alpha_{i})\right]\,\,.$ (2)
This regularization promotes binary opacity distributions, allowing low-opacity Gaussians to be more aggressively pruned during densification. Incorporating this term significantly sharpens geometric reconstruction, providing a better foundation for subsequent synthesis.
3.1.4 Pseudo camera depth supervision.
To further reduce floating artifacts, we sample pseudo-cameras positioned closer to the ground during optimization. From these pseudo-cameras, we render RGB images $I_{\text{RGB}}$ and corresponding alpha-blended depth maps $\hat{D}_{\text{GS}}$. We then use an off-the-shelf monocular depth estimator, MoGe [90], to predict scale-invariant depths $\hat{D}_{\text{est}}$ from these renders. We use the absolute value of Pearson correlation (PCorr) to supervise the depth:
$\small\mathcal{L}_{\text{depth}}=1-\lVert\text{PCorr}(\hat{D}_{\text{GS}},\hat{D}_{\text{est}})\rVert_{1}\,\,;\quad\text{PCorr}(\hat{D}_{\text{GS}},\hat{D}_{\text{est}})=\frac{\text{Cov}(\hat{D}_{\text{GS}},\hat{D}_{\text{est}})}{\sqrt{\text{Var}(\hat{D}_{\text{GS}})\text{Var}(\hat{D}_{\text{est}})}}\,\,.$ (3)
3.1.5 Optimization.
Combining all components, the overall loss for the reconstruction stage is defined as:
$\small\mathcal{L}_{\text{sat}}(G,C)=\mathcal{L}_{\text{color}}+\lambda_{\text{op}}\mathcal{L}_{\text{op}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}\,\,$, (4)
@@
3.2 Synthesize via Curriculum-Based Iterative Datasets Update
The iterative dataset update (IDU) technique [23, 59] repeatedly executes render-edit-update cycles across multiple episodes to progressively synthesize 3D scenes. Unlike previous methods that sample camera poses from original training views [23] or simple orbits [59], we introduce a curriculum-based refinement schedule over $N_{e}$ episodes that specifically addresses the geometric and visual limitations inherent to satellite imagery, producing structurally accurate and photorealistic reconstructions of occluded areas.
3.2.1 Curriculum learning strategy.
As illustrated in Figure˜5, we observe that 3DGS trained from satellite imagery produces higher-quality renders at higher elevation angles but degenerates at lower elevation angles. Leveraging this insight, we introduce a curriculum-based synthesis strategy, which progressively lowers viewpoints across optimization episodes. Specifically, we define $N_{p}$ look-at points $\{P_{j}\}_{j=1}^{N_{p}}$ uniformly placed throughout the scene and uniformly sample $N_{v}$ camera positions along orbital trajectories with controlled elevation angles and radii. Our iterative dataset update (IDU) process starts at higher elevations and progressively descends toward lower perspectives. This approach gradually reveals previously occluded regions, improving geometric detail and texture realism, as validated in our ablation studies (Section˜4.2).
3.2.2 Render refinement by text-to-image diffusion model.
As illustrated in Figure˜5(a), renderings from initial 3DGS contain blurry texture and artifacts. To address this, we leverage prompt-to-prompt editing with a pre-trained text-to-image diffusion model to synthesize disocclusion areas, remove artifacts, and enhance geometry. Prompt-to-prompt editing [24] modifies input images, which are described by the source prompt, to align with the target prompt while preserving structural content. Although originally designed for real or synthetically generated images, we demonstrate its effectiveness for refining degraded 3DGS renders produced from satellite-view training, enabling high-quality appearance enhancement without disrupting the scene’s satellite-consistent geometry. Specifically, we use FlowEdit [36] with the FLUX.1 [dev] diffusion model [39]. Our source prompts describe the original degraded features. The target prompts specify the desired high-quality attributes (details in Supplementary). This approach significantly improves rendering quality. It yields sharper geometry, richer textures, and physically coherent shadows (see Figure˜5). This strengthens the 3DGS training dataset for more accurate reconstructions.
3.2.3 Multiple diffusion samples.